Solve by factoring
3x^2+5x-2=0

Answer by shree840(260)

3x^2+5x-2=0
you multiply 3 with -2=-6
get possible factors 1*6 2*3
choose so that when you multiply you get -6 and when you add you get +5
3x^2+6x-1x-2=0
3x(x+2)-1(x+2)=0
(3x-1)(x+2)=0 so either 3x-1=0 or x=1/3
or x+2=0 or x=-2